# Optimal energy storage and collective charging speedup in the central-spin quantum battery

### v1

@@ Abstract
Quantum batteries (QBs) exploit principles of quantum mechanics to accelerate the charging process and aim to achieve optimal energy storage. However, analytical results for investigating these problems remain lacking due to the challenges associated with nonequilibrium dynamics. In this work, we analytically investigate a central-spin QB model in which $N_{b}$ spin-1/2 battery cells interact with $N_{c}$ spin-1/2 charger units, using $m$ initially excited charger units as a resource. By employing the invariant subspace method and the shifted Holstein-Primakoff (HP) transformation, we identify four scenarios in which optimal energy storage can be achieved: 1) $N_{b}\!\ll\!m\!\ll\!N_{c}$; 2) $m\!\ll\!N_{b}\!\ll\!N_{c}$; 3) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and 4) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$]. In these cases, optimal storage is ensured by the SU(2) symmetry emerging from the charging dynamics. The first three cases map the central-spin QB to different Tavis-Cummings (TC) QBs, while the fourth corresponds to the non-TC limit. We analytically determine the charging time and demonstrate that in the fully charging cases (1) and (4), the collective charging exhibits an $N_{b}$-fold enhancement in speedup compared to the parallel charging scheme. Additionally, we numerically observe a unified charging behavior when $m\!=\!N_{c}$, showing that asymptotically optimal energy storage is possible when $N_{b}\!=\!m\!=\!N_{c}$. In this case, we find a collective charging enhancement scaling as $N_{b}^{0.8264}$. Our results highlight the crucial role of dynamically emergent SU(2) symmetry in providing an analytical understanding of non-equilibrium charging dynamics in QBs.
Figure 1: Schematic diagram of the central-spin QB. (a) Initially, the system consists of $N_{b}$ battery cells in their ground states, with no excitations, and $N_{c}$ charger units prepared in a Dicke state containing $m$ excitations. Comparison between (b) the collective charging scheme and (c) the parallel charging scheme.
@@ I Introduction
Pioneering theoretical work on energy extraction from many-body states has suggested that the generation of multipartite entanglement can enhance the efficiency of work extraction Hovhannisyan et al. (2013). This finding implies the existence of a new class of energy storage devices, known as quantum batteries Alicki and Fannes (2013) (QBs), which exploit unique quantum features to accelerate the charging process and outperform classical batteries. Compared to charging individual battery cells independently, collective charging exhibits a significant quadratic speedup in charging time Gyhm et al. (2022); Campaioli et al. (2017); Julià-Farré et al. (2020), which is believed to arise from multipartite entanglement induced by collective interactions. The phenomenon where charging power scales faster than the number of quantum cells is referred to as collective charging advantage Campaioli et al. (2024), which has been observed in various models, including central-spin model Peng et al. (2021), spin-chain model Le et al. (2018); Huangfu and Jing (2021), Sachdev–Ye–Kitaev model Rossini et al. (2020) and Dicke model Zhang and Blaauboer (2023); Ferraro et al. (2018); Dou et al. (2022).
@@
On the other hand, energy storage is another crucial metric for evaluating the performance of a quantum battery. Practically, a QB that charges rapidly but stores only minimal energy is not desirable. Recent studies have explored the relationship between energy storage and quantum correlations Shi et al. (2022); Andolina et al. (2019); Šafránek et al. (2023). However, identifying many-body models that satisfy the conditions for achieving optimal energy storage remains challenging. Here, optimal energy storage refers to a situation where the battery absorbs all the energy initially stored in the charger or reaches a fully charged state, which will be rigorously defined later. Our recent work has shown that the emergence of SU(2) symmetry in the Tavis-Cummings (TC) model guarantees the realization of optimal energy storage Yang et al. (2024). Could this emergent symmetry also explain the presence of optimal energy storage in other models? Moreover, under conditions of optimal energy storage, can we achieve collective charging speedup simultaneously? We aim to address these questions in an important low-dimensional model.
@@
The central-spin model has provided significant insights into the quantitative understanding of decoherence in nitrogen-vacancy centers in diamond, as well as the dynamics of entanglement Schliemann et al. (2003); Quan et al. (2006); Hanson et al. (2008); Bortz et al. (2010); Barnes et al. (2012); Lu et al. (2020); Wan et al. (2020); Claeys et al. (2018); Wu et al. (2020); He et al. (2019); Stanek et al. (2013); Cucchietti et al. (2005); Lee et al. (2008); Onizhuk et al. (2024); Rossini et al. (2007). This model can be naturally put into the framework of QBs, referred to as the central-spin QB, where $N_{b}$ central spins act as battery cells and $N_{c}$ bath spins serve as charging units. It has been argued that a collective charging advantage is present in this model; however, the analysis is based on applying the Holstein-Primakoff (HP) transformation to both the bath and central spins Peng et al. (2021), which is valid only in the short-time regime and does not extend to the time scale including maximal energy storage Andolina et al. (2019). On the other hand, the use of the HP approximation simplifies the central-spin model to the TC model, which we refer to as the TC limit. As discussed in Ref. Yang et al. (2024), even within the TC model, optimal energy storage is not guaranteed in all cases. In a specific non-TC limit Liu et al. (2021), we identified an optimal energy storage scenario for the case of $N_{b}\!=\!2$ and conjectured, based on numerical evidence, that optimal energy storage persists under this condition for arbitrary $N_{b}$ case Liu et al. (2021). The corresponding charging time has also been conjectured and numerically validated. Despite these results, analytical results regarding the conditions for optimal energy storage and the determination of charging time in the central-spin model remain lacking. In particular, can we provide rigorous proof for the conditions and charging time for optimal energy storage in the non-TC limit? Additionally, even for the TC limit, different initial configurations in the central-spin model can be used to realize it. Do these various configurations all guarantee optimal energy storage? Do they yield the same charging time? Addressing these questions will help distinguish the central-spin model from its behavior under the TC limit and deepen our understanding of the nonequilibrium dynamics in the central-spin model.
@@
In this paper, we address these questions systematically. In Sec. II, we introduce the central-spin model and the invariant subspace approach used for the subsequent analysis. The quantities that characterize optimal energy storage and the collective charging advantage are also given. Section III employs the HP transformation to identify four initial configurations that achieve the TC limits. For the first three cases, we rigorously demonstrate the emergent SU(2) symmetry that ensures optimal energy storage and analytically determine the corresponding charging time. Based on these findings, we prove that the collective charging advantage scales with $N_{b}$ in the first fully charged TC limit. For the fourth case, numerical calculations indicate the absence of optimal energy storage. In Sec. IV, we explore the non-TC limits. We analytically demonstrate a novel scenario for achieving optimal energy storage while maintaining an $N_{b}$ scaling for the collective charging advantage. Additionally, numerical analysis reveals another non-TC configuration that attains asymptotic optimal energy storage, with a collective charging advantage scaling as $N_{b}^{0.8264}$. Finally, a conclusion is made in Sec. V.
@@ II Central-spin quantum battery
The central-spin QB is described by the following Hamiltonian
@@
Initially, the battery is set to the ground state of $H_{b}$ with $N_{b}$ spins down, namely, $|0\rangle_{b}\!\equiv|\!\downarrow_{1},\downarrow_{2},\cdot\cdot\cdot,\downarrow_{N_{b}}\rangle_{b}$, while the charger is prepared in a high-energy Dicke state $|m\rangle_{c}$ with $m$ spins up, see Fig.1(a). The charging dynamics are performed by turning on the interactions $H_{I}$. To ensure that the energy injected into the battery exclusively comes from the charger, we require $[H_{b}+H_{c},H]\!=\!0$ and thus we only consider the resonant case with $\omega_{b}\!=\!\omega_{c}\!\equiv\!\omega$. Therefore, the free part in the Hamiltonian (II) is a constant and we will ignore it. In terms of the invariant subspace $\mathcal{H}_{m}=\{|0\rangle_{b}|m\rangle_{c},|1\rangle_{b}|m\!-\!1\rangle_{c},\cdot\cdot\cdot,|d\rangle_{b}|m\!-\!d\rangle_{c}\}$ containing the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$, the Hamiltonian $H$ defined in Eq. (II) can be represented as a $(d+1)\times(d+1)$ matrix in the invariant subspace $\mathcal{H}_{m}$ by using Dicke basis
@@
where we have shifted the ground state energy to cancel the diagonal terms. The off-diagonal terms are given by
@@
The dimension of the effective Hamiltonian (2) is $(d\!+\!1)$ with $d\equiv\min\{{N_{b},m}\}$, which scales linearly with the minimal number of battery cells or the charger units. Therefore, it is easy to diagonalize $\bm{H}$ for calculating dynamics $|\psi(t)\rangle=e^{-iHt}|\psi(0)\rangle$, details can be found in Ref. Liu et al. (2021).
@@
which depends solely on the battery size $N_{b}$, the charger size $N_{c}$, and the initial number of excited charger spins $m$. For a fair comparison, the resources used in the parallel charging scheme [see Fig. 1(c)] must be equivalent to those used in the collective charging scheme. In the parallel charging scheme, there are $N_{b}$ individual single-cell QBs, where each battery cell interacts with $N_{c}/N_{b}$ charger units, with $m/N_{b}$ spins initially excited. To make sense for this setting, we require that $N_{c}\!\geq\!N_{b}$ and $m\!\geq\!N_{b}$, and we will define the collective charging advantage Eq. (9) only for this case. According to Ref. Liu et al. (2021), the charging power for a single-cell central-spin QB is given by
@@
Figure 2: Charging dynamics for four TC limits: a) $N_{b}\!\ll\!m\!\ll\!N_{c}$; b) $m\!\ll\!N_{b}\!\ll\!N_{c}$; c) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and d) $m\!=\!N_{c}\!\ll\!N_{b}$. The blue lines are numerically obtained and the red lines are theoretical results. The coupling strength is set to $A\!=\!1$.
@@ III Optimal energy storage and collective charging speedup at TC Limits
In previous work Liu et al. (2021), we analytically demonstrated a case of optimal energy storage for $2\!=\!N_{b}\!\ll\!m\!\ll\!N_{c}$ and numerically verified the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$ with arbitrary $N_{b}$ to achieve optimal energy storage. However, the corresponding charging time was not determined. Actually, under these initial conditions, the central-spin model can be mapped to the TC model, which we refer to as the TC limit. In this section, we will extend our previous work to find more TC limits and then we will use the theory developed in Ref. Yang et al. (2024) to determine not only the possibility for achieving optimal energy storage but also the collective charging advantage. We emphasize that the emergent SU(2) symmetry in the dynamics plays a crucial role in achieving optimal energy storage and is key to analytically calculating the collective charging advantage.
For the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$, due to the conservation of the total excitation number $[J^{z}\!+\!S^{z},H]\!=\!0$, the number of excitations in the charger part cannot exceed the initial value $m$. Since $\lim_{N_{c}\to\infty}m/N_{c}=0$, the charger remains close to the vacuum state $|0\rangle_{c}$ throughout the dynamics at the limit $N_{c}\!\to\!\infty$ Therefore, we can apply the HP transformation to the charger Hamiltonian as follows
@@
where $a$ and $a^{\dagger}$ are bosonic annihilation and creation operators. In this correspondence, the number of charger spin excitations $m$ in the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$ corresponds to the photon number. In Ref. Yang et al. (2024), we have shown that for TC battery (11) and initial state $|\psi_{\rm TC}(0)\rangle=|0\rangle_{b}|n_{0}\rangle_{c}$ with $|n_{0}\rangle$ being a Fock state with $n_{0}$ photons, a distinct SU(2) symmetry emerges from the dynamics for the limiting cases $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$. This results in optimal energy storage given by $\Delta E(T_{\rm TC})\!=\!\omega\min\{N_{b},n_{0}\}$ with the charging time
@@
Therefore, we conclude that optimal energy storage in the central-spin QB can be achieved with the initial state $|\psi(0)\rangle=|0\rangle_{b}|m\rangle_{c}$ with $N_{b}\!\ll\!m\!\ll\!N_{c}$, see Fig.2(a). By substituting $n_{0}\!=\!m$ into Eq. (12), the charging time is given by
@@
which is numerically verified in Fig. 2(a). Then by substituting Eqs. (7) and (13) into Eq. (9), the collective charging advantage is obtained as follows
@@
It is important to emphasize that the TC Hamiltonian (11), which describes the central-spin QB, only requires the number of excited charger spins to be much smaller than the total number of charger spins, i.e., $m\!\ll\!N_{c}$. Moreover, the condition $n_{0}\!\ll\!N_{b}$ for optimal energy storage in the TC battery implies another scenario: $m\!\ll\!N_{b}\!\ll\!N_{c}$, which also facilitates optimal energy storage in the central-spin QB, see Fig. 2(b). By setting $n_{0}\!=\!m$ in Eq. (12), we obtain the charging time as follows
@@
Unlike the two situations discussed above, we can also employ the HP transformation to map the battery onto a bosonic mode, facilitating the construction of a TC QB from the central-spin QB. Now, let us consider the case where $m\!\ll\!N_{c}\!\ll\!N_{b}$. In this regime, the possible maximum excitation for the battery part is $m$, which is significantly smaller than the battery size $N_{b}$. Consequently, the HP transformation can be applied to the battery spins, yielding
@@
where the battery has become a photon field. From the perspective of matrix representation in the invariant subspace, Eq. (2), the above HP transformation approximates the matrix element $u_{j}$ as
@@
where $\Omega\equiv 2A\sqrt{N_{c}-(m-1)/2}$ is the generalized Rabi frequency and $\bm{L^{x}}$ is $x$-direction generator of SU(2) algebra in the spin-$m$ representation. The evolution of energy storage (4) can be represented as
@@
$=\frac{\pi}{2A\sqrt{N_{b}}\sqrt{N_{c}-\frac{m-1}{2}}},$ (24)
see Fig.2(c) for numerical verification.
@@
Figure 3: (a) Charging dynamics in the non-TC limit: $N_{b}\!\ll\!m\!=\!kN_{c}$ with $10\!<\!k\!<\!$ 1. The blue lines represent numerical results, while the red lines correspond to theoretical predictions. (b) The analytical charging time, Eq. (33), as a function of $k$ is numerically verified. The coupling strength is set to $A\!=\!1$.

### v2

@@ Abstract
Quantum batteries (QBs) exploit principles of quantum mechanics to accelerate the charging process and aim to achieve optimal energy storage. However, analytical results for investigating these problems remain lacking due to the challenges associated with nonequilibrium dynamics. In this work, we analytically investigate a central-spin QB model in which $N_{b}$ spin-1/2 battery cells interact with $N_{c}$ spin-1/2 charger units, using $m$ initially excited charger units as a resource. By employing the invariant subspace method and the shifted Holstein-Primakoff transformation, we identify four scenarios in which optimal energy storage can be achieved: (i) $N_{b}\!\ll\!m\!\ll\!N_{c}$; (ii) $m\!\ll\!N_{b}\!\ll\!N_{c}$; (iii) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and (iv) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$]. In these cases, optimal storage is ensured by the SU(2) symmetry emerging from the charging dynamics. The first three cases map the central-spin QB to different Tavis-Cummings (TC) QBs, while the fourth corresponds to the non-TC limit. We analytically determine the charging time and demonstrate that in the fully charging cases (i) and (iv), the collective charging exhibits an $N_{b}$-fold enhancement in speedup compared to the parallel charging scheme. Additionally, we numerically observe a unified charging behavior when $m\!=\!N_{c}$, showing that asymptotically optimal energy storage is possible when $N_{b}\!=\!m\!=\!N_{c}$. In this case, we find a collective charging enhancement scaling as $N_{b}^{0.8264}$. The origin of the collective charging advantage in central-spin quantum batteries is also analyzed through the quantum speed limit and a multipartite entanglement witness. Our results highlight the crucial role of dynamically emergent SU(2) symmetry in providing an analytical understanding of non-equilibrium charging dynamics in QBs.
Figure 1: Schematic diagram of the central-spin QB. (a) Initially, the system consists of $N_{b}$ battery cells in their ground states, with no excitations, and $N_{c}$ charger units prepared in a Dicke state containing $m$ excitations. Also shown is a comparison between (b) the collective charging scheme and (c) the parallel charging scheme.
@@ I Introduction
Pioneering theoretical work on energy extraction from many-body states has suggested that the generation of multipartite entanglement can enhance the efficiency of work extraction [1]. This finding implies the existence of a new class of energy storage devices, known as quantum batteries [2] (QBs), which exploit unique quantum features to accelerate the charging process and outperform classical batteries. Compared to charging individual battery cells independently, collective charging exhibits a significant quadratic speedup in charging time [3, 4, 5, 6], which is believed to arise from multipartite entanglement induced by collective interactions. The phenomenon where charging power scales faster than the number of quantum cells is referred to as collective charging advantage [7], which has been observed in various models, including the central-spin model [8], spin-chain model [9, 10], Sachdev–Ye–Kitaev model [11], and Dicke model [12, 13, 14].
@@
On the other hand, energy storage is another crucial metric for evaluating the performance of a quantum battery [15, 16, 17]. Practically, a QB that charges rapidly but stores only minimal energy is not desirable. Recent studies have explored the relationship between energy storage and quantum correlations [15, 18, 19]. However, identifying many-body models that satisfy the conditions for achieving optimal energy storage remains challenging. Here optimal energy storage refers to a situation where the battery absorbs all the energy initially stored in the charger or reaches a fully charged state, which will be rigorously defined later. Our recent work has shown that the emergence of SU(2) symmetry in the Tavis-Cummings (TC) model guarantees the realization of optimal energy storage [16]. Could this emergent symmetry also explain the presence of optimal energy storage in other models? Moreover, under conditions of optimal energy storage, can we achieve collective charging speedup simultaneously? We aim to address these questions in an important low-dimensional model.
@@
The central-spin model has provided significant insights into the quantitative understanding of decoherence in nitrogen-vacancy (NV) centers in diamond, as well as the dynamics of entanglement [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. This model can be naturally put into the framework of QBs, referred to as the central-spin QB, where $N_{b}$ central spins act as battery cells and $N_{c}$ bath spins serve as charging units. It has been argued that a collective charging advantage is present in this model; however, the analysis is based on applying the Holstein-Primakoff (HP) transformation to both the bath and central spins [8], which is valid only in the short-time regime and does not extend to the time scale including maximal energy storage [18]. On the other hand, the use of the HP approximation simplifies the central-spin model to the TC model, which we refer to as the TC limit. As discussed in Ref. [16], even within the TC model, optimal energy storage is not guaranteed in all cases. In a specific non-TC limit [35], we identified an optimal energy storage scenario for the case of $N_{b}\!=\!2$ and conjectured, based on numerical evidence, that optimal energy storage persists under this condition for the arbitrary $N_{b}$ case [35]. The corresponding charging time has also been conjectured and numerically validated. Despite these results, analytical results regarding the conditions for optimal energy storage and the determination of charging time in the central-spin model remain lacking. In particular, can we provide rigorous proof for the conditions and charging time for optimal energy storage in the non-TC limit? Additionally, even for the TC limit, different initial configurations in the central-spin model can be used to realize it. Do these various configurations all guarantee optimal energy storage? Do they yield the same charging time? Addressing these questions will help distinguish the central-spin model from its behavior under the TC limit and deepen our understanding of the nonequilibrium dynamics in the central-spin model.
@@
In this paper we address these questions systematically. In Sec. II we introduce the central-spin model and the invariant subspace approach used for the subsequent analysis. The quantities that characterize optimal energy storage and the collective charging advantage are also given. Section III employs the HP transformation to identify four initial configurations that achieve the TC limits. For the first three cases, we rigorously demonstrate the emergent SU(2) symmetry that ensures optimal energy storage and analytically determine the corresponding charging time. Based on these findings, we prove that the collective charging advantage scales with $N_{b}$ in the first fully charged TC limit. For the fourth case, numerical calculations indicate the absence of optimal energy storage. In Sec. IV we explore the non-TC limits. We analytically demonstrate a novel scenario for achieving optimal energy storage while maintaining an $N_{b}$ scaling for the collective charging advantage. Additionally, numerical analysis reveals another non-TC configuration that attains asymptotic optimal energy storage, with a collective charging advantage scaling as $N_{b}^{0.8264}$. The origin of the collective charging advantage is discussed in Sec. V using the quantum speed limit and a multipartite entanglement witness. A summary is given in Sec. VI.
@@ II Central-spin quantum battery
The central-spin QB is described by the Hamiltonian
@@
Initially, the battery is set to the ground state of $H_{b}$ with $N_{b}$ spins down, namely, $|0\rangle_{b}\!\equiv|\!\downarrow_{1},\downarrow_{2},\cdot\cdot\cdot,\downarrow_{N_{b}}\rangle_{b}$, while the charger is prepared in a high-energy Dicke state $|m\rangle_{c}$ with $m$ spins up [see Fig.1(a)]. The charging dynamics is performed by turning on the interactions $H_{I}$. To ensure that the energy injected into the battery exclusively comes from the charger, we require $[H_{b}+H_{c},H]\!=\!0$ and thus we only consider the resonant case with $\omega_{b}\!=\!\omega_{c}\!\equiv\!\omega$. Therefore, the free part in the Hamiltonian (II) is a constant and we will ignore it. In terms of the invariant subspace $\mathcal{H}_{m}=\{|0\rangle_{b}|m\rangle_{c},|1\rangle_{b}|m\!-\!1\rangle_{c},\cdot\cdot\cdot,|d\rangle_{b}|m\!-\!d\rangle_{c}\}$ containing the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$, the Hamiltonian $H$ defined in Eq. (II) can be represented as a $(d+1)\times(d+1)$ matrix in the invariant subspace $\mathcal{H}_{m}$ by using the Dicke basis
@@
where we have shifted the ground-state energy to cancel the diagonal terms. The off-diagonal terms are given by
@@
The dimension of the effective Hamiltonian (2) is $d\!+\!1$ with $d\equiv\min\{{N_{b},m}\}$, which scales linearly with the minimal number of battery cells or the charger units. Then it is easy to diagonalize $\bm{H}$ to calculate the dynamics
[…]
Thus, the evolution of the total quantum state is expressed as
$|\psi(t)\rangle=\bm{\psi}_{0}(t)|0\rangle_{b}|m\rangle_{c}+\cdots+\bm{\psi}_{d}(t)|d\rangle_{b}|m-d\rangle_{c}.$
The reduced density matrix of the battery state is then obtained as
$\rho_{b}(t)$$\equiv$ $\mathrm{Tr}_{c}[|\psi(t)\rangle\langle\psi(t)|]$ (6)
$=$ $|\bm{\psi}_{0}(t)|^{2}|0\rangle\langle 0|+\cdots+|\bm{\psi}_{d}(t)|^{2}|d\rangle\langle d|.$
@@
which depends solely on the battery size $N_{b}$, the charger size $N_{c}$, and the initial number of excited charger spins $m$. For a fair comparison, the resources used in the parallel charging scheme [see Fig. 1(c)] must be equivalent to those used in the collective charging scheme. In the parallel charging scheme, there are $N_{b}$ individual single-cell QBs, where each battery cell interacts with $N_{c}/N_{b}$ charger units, with $m/N_{b}$ spins initially excited. To make sense for this setting, we require that $N_{c}\!\geq\!N_{b}$ and $m\!\geq\!N_{b}$, and we will define the collective charging advantage (12) only for this case. According to Ref. [35], the charging power for a single-cell central-spin QB is given by
@@
Figure 2: Charging dynamics for four TC limits: (a) $N_{b}\!\ll\!m\!\ll\!N_{c}$; (b) $m\!\ll\!N_{b}\!\ll\!N_{c}$; (c) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and (d) $m\!=\!N_{c}\!\ll\!N_{b}$. The blue lines are numerically obtained and the red lines are theoretical results. The coupling strength is set to $A\!=\!1$.
@@ III Optimal energy storage and collective charging speedup at TC Limits
In previous work [35] we analytically demonstrated a case of optimal energy storage for $2\!=\!N_{b}\!\ll\!m\!\ll\!N_{c}$ and numerically verified the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$ with arbitrary $N_{b}$ to achieve optimal energy storage. However, the corresponding charging time was not determined. Actually, under these initial conditions, the central-spin model can be mapped to the TC model, which we refer to as the TC limit. In this section we extend our previous work to find more TC limits and then we use the theory developed in Ref. [16] to determine not only the possibility for achieving optimal energy storage but also the collective charging advantage. We emphasize that the emergent SU(2) symmetry in the dynamics plays a crucial role in achieving optimal energy storage and is key to analytically calculating the collective charging advantage.
For the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$, due to the conservation of the total excitation number $[J^{z}\!+\!S^{z},H]\!=\!0$, the number of excitations in the charger part cannot exceed the initial value $m$. Since $\lim_{N_{c}\to\infty}m/N_{c}=0$, the charger remains close to the vacuum state $|0\rangle_{c}$ throughout the dynamics at the limit $N_{c}\!\to\!\infty$ Therefore, we can apply the HP transformation to the charger Hamiltonian as
@@
where $a$ and $a^{\dagger}$ are bosonic annihilation and creation operators, respectively. In this correspondence, the number of charger spin excitations $m$ in the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$ corresponds to the photon number. In Ref. [16] we have shown that for TC battery (14) and initial state $|\psi_{\rm TC}(0)\rangle=|0\rangle_{b}|n_{0}\rangle_{c}$, with $|n_{0}\rangle$ being a Fock state with $n_{0}$ photons, a distinct SU(2) symmetry emerges from the dynamics for the limiting cases $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$. This results in optimal energy storage given by $\Delta E(T_{\rm TC})\!=\!\omega\min\{N_{b},n_{0}\}$ with the charging time
@@
Therefore, we conclude that optimal energy storage in the central-spin QB can be achieved with the initial state $|\psi(0)\rangle=|0\rangle_{b}|m\rangle_{c}$, with $N_{b}\!\ll\!m\!\ll\!N_{c}$ [see Fig.2(a)]. By substituting $n_{0}\!=\!m$ into Eq. (15), the charging time is given by
@@
which is numerically verified in Fig. 2(a). Then by substituting Eqs. (10) and (16) into Eq. (12), the collective charging advantage is obtained as
@@
It is important to emphasize that the TC Hamiltonian (14), which describes the central-spin QB, only requires the number of excited charger spins to be much smaller than the total number of charger spins, i.e., $m\!\ll\!N_{c}$. Moreover, the condition $n_{0}\!\ll\!N_{b}$ for optimal energy storage in the TC battery implies another scenario $m\!\ll\!N_{b}\!\ll\!N_{c}$, which also facilitates optimal energy storage in the central-spin QB [see Fig. 2(b)]. By setting $n_{0}\!=\!m$ in Eq. (15), we obtain the charging time as
@@
Unlike the two situations discussed above, we can also employ the HP transformation to map the battery onto a bosonic mode, facilitating the construction of a TC QB from the central-spin QB. Now let us consider the case where $m\!\ll\!N_{c}\!\ll\!N_{b}$. In this regime, the possible maximum excitation for the battery part is $m$, which is significantly smaller than the battery size $N_{b}$. Consequently, the HP transformation can be applied to the battery spins, yielding
@@
where the battery has become a photon field. From the perspective of matrix representation in the invariant subspace [Eq. (2)], the above HP transformation approximates the matrix element $u_{j}$ as
@@
where $\Omega\equiv 2A\sqrt{N_{c}-(m-1)/2}$ is the generalized Rabi frequency and $\bm{L^{x}}$ is the $x$-direction generator of SU(2) algebra in the spin-$m$ representation. The evolution of energy storage (7) can be represented as
@@
$=\frac{\pi}{2A\sqrt{N_{b}}\sqrt{N_{c}-\frac{m-1}{2}}}$ (27)
[see Fig.2(c) for numerical verification].
@@
Figure 3: (a) Charging dynamics in the non-TC limit: $N_{b}\!\ll\!m\!=\!kN_{c}$ with $10\!<\!k\!<\!$ 1. The blue lines represent numerical results, while the red lines correspond to theoretical predictions. (b) The analytical charging time (36) as a function of $k$ is numerically verified for $N_{b}\!\ll\!m\!=\!kN_{c}$ with $10\!<\!k\!<\!$ 1. The coupling strength is set to $A\!=\!1$.